OpenStudy (jmartinez638):

Calculate dy/dx if 2cos^2 *y=xy

OpenStudy (jmartinez638):

Calculate dy/dx if \[2\cos^2y=xy\]

OpenStudy (solomonzelman):

differentiate each side (note that the derivative of any "y" will have a chain rule of y' being that y is also a function of x)

OpenStudy (solomonzelman):

What would be the derivative of \(\cos^2x\) ? Once you calculate that, the derivative of \(\cos^2y\) will be the same, EXCEPT that it will be multiplied times y'.

OpenStudy (jmartinez638):

so it would be -2cos(x)sin(x)=y ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2\cos^2x }\) would have a derivative of \(\large\color{black}{ \displaystyle -2\cos(x)\sin(x) }\) AND, \(\large\color{black}{ \displaystyle 2\cos^2y }\) would have a derivative of \(\large\color{black}{ \displaystyle -2\cos(y)\sin(y) \times y'}\)

OpenStudy (solomonzelman):

this y', is simply a chain rule.

OpenStudy (solomonzelman):

then the right side, do you know how to differentiate that?

OpenStudy (jmartinez638):

It would end up either x^0y=y*1 or dy/dx[xy] so that'd be dy/d[y]? or something like that

OpenStudy (solomonzelman):

the derivative of xy goes by the product rule: \(\large\color{black}{ \displaystyle \frac{dy}{dx} \left[ xy \right]=x\frac{dy}{dx} \left[ y \right]+y\frac{dy}{dx} \left[ x \right] =xy'+y\cdot1}\)

OpenStudy (jmartinez638):

So what would this derivative look like in final form? not too much simplification would be required, luckily.

OpenStudy (solomonzelman):

yes, just: \(\large\color{black}{ \displaystyle 2\cos^2 y=xy }\) differentiate: \(\large\color{black}{ \displaystyle -2y'\sin y\cos y=xy'+y }\) solve for y' \(\large\color{black}{ \displaystyle -2y'\sin y\cos y-xy'=y }\) \(\large\color{black}{ \displaystyle y'(-2\sin y\cos y-x)=y }\) \(\large\color{black}{ \displaystyle y'=y/(-2\sin y\cos y-x) }\)

OpenStudy (zarkon):

\[\large\color{black}{ \displaystyle -\color{red}{4}y'\sin y\cos y=xy'+y }\]

OpenStudy (solomonzelman):

oh, yeah, there is another 2, thanks for correcting me

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2\cos^2 y=xy }\) differentiate: \(\large\color{black}{ \displaystyle -4y'\sin y\cos y=xy'+y }\) solve for y' \(\large\color{black}{ \displaystyle -4y'\sin y\cos y-xy'=y }\) \(\large\color{black}{ \displaystyle y'(-4\sin y\cos y-x)=y }\) \(\large\color{black}{ \displaystyle y'=y/(-4\sin y\cos y-x) }\)

OpenStudy (jmartinez638):

Thanks, that really helped me out!